The equation for an alternating current is given by i - 77 sin 314 t. Find the peak value, frequency, time period and instantaneous value at t-2 ms
1
Expert's answer
2020-11-10T07:02:20-0500

Let us write the equation for alternating current in general form

"i(t)=I_m \cdot Sin{(2 \pi f \cdot t)}"

Then

Peak current "I_m=77A"

Frequency "f=\frac{314}{2\pi}=\frac{314}{2\cdot 3.14}=50 \frac{1}{s}"

Time period "T=\frac{1}{f}=\frac{1}{50}=0.02s"

Instantaneous value at t=2 ms.

"i(0.002)=77 \cdot Sin{(314 \cdot 0.002)}=45.24A"